# LLM-Oracle Machines

### v2

@@ 2.2 Augmented LLM-OM
Ideally, the answer $A$ to the inquiry $Q$ should align with the augmented text $T$. Here’s a formal definition of this alignment.
  • Let $\mbox{mng}(X)$ denote the set of meanings of text $X$. We say that answer $A$ complies with augmented text $T$ with respect to the inquiry $Q$, written as $A\lesssim_{Q}T$, if these two conditions hold:
@@
      Answer $A$ is relevant to the inquiry $Q$. Namely, $Q\lesssim A$.
@@
      The meaning of answer $A$ is a subset of the meaning of augmented text $T$. That is, $\mbox{mng}(A)\subseteq\mbox{mng}(T)$.
@@ 2.3 Fault-avoidance LLM-OM
  • $M$ is considered consistent if, for any user inquiries $Q$ and $Q^{\prime}$ posed at different points in time, if $\mbox{mng}(Q^{\prime})=\mbox{mng}(Q)$, then $M$ outputs an answer $A$ to $Q$ and an answer $A^{\prime}$ to $Q^{\prime}$ such that $\mbox{mng}(A)=\mbox{mng}(A^{\prime})$.
  • Suppose $M$ is augmented with input and output being $(T,Q)$ and $A$. We say that $A$ is correct for $Q$ with respect to $T$ if the following three conditions hold:
@@
      $A$ complies with the ground truth $T$ with respect to $Q$. Namely, $A\lesssim_{Q}T$.
@@
      $A$ doesn’t contain information not expressed in $Q$. Namely, $\mbox{mng}(\overline{Q})\cap\mbox{mng}(A)=\emptyset$, where $\mbox{mng}(\overline{Q})$ represents any other meanings not expressed in $Q$.
@@
      $A$ is complete for $Q$ with respect to $T$. Namely, $A$ doesn’t miss any information contained in $T$ that are relevant to $Q$. In other words, $\mbox{mng}(Q)\cap(\mbox{mng}(T)-\mbox{mng}(A))=\emptyset$.
    Then $M$ is correct with respect to $T$ if for any input $(T,Q)$, $M$ returns an answer $A$ that is correct for $Q$ with respect to $T$.
@@
      Including Additional Information: Answer $A$ may contain information that extend beyond the specific information requested in the user inquiry $Q$.
@@
      Incompleteness: $A$ might be incomplete for $Q$ with respect to $T$. Namely, it doesn’t capture all the relevant details from the augmented text $T$ that pertain to $Q$.
@@
      Distributional discrepancy: The distribution of meanings within answer $A$, denoted as $\mbox{dist}(A)$, might be different from the distribution of meanings in the relevant portion of the augmented text $T$ with respect to $A$, denoted as $\mbox{dist}(T_{A})$.
    Essentially, adequacy acknowledges that the answer $A$ may not be perfect, but it still provides some valuable insights related to the query and the augmented text.
@@
    Then $M$ is adequate if for all inputs $(T,Q)$, $M$ returns an answer $A$ that is adequate for $Q$ with respect to $T$.
@@
Verifying the correctness or adequacy of the answer $A$ for a given query $Q$ with respect to $T$ calls for concrete implementations of $\mbox{toc}(X)$ and $\mbox{mng}(X)$ using techniques in natural language processing, machine learning, deep learning, and other methods.
Consistency, however, is much more difficult to verify. We may aim to develop a method that provides a certain guarantee that $M$ is consistent with a desired high probability.

### v3

@@ 2.2 Augmented LLM-OM
Ideally, the answer $A$ to the inquiry $Q$ should align with the augmented text $T$. Here’s a formal definition of this alignment. Let $\mbox{mng}(X)$ denote the set of meanings of text $X$.
  • We say that $A$ complies with $T$ with respect to $Q$, written as $A\lesssim_{Q}T$, if the following conditions hold:
@@
      Relevance: $A$ is relevant to $Q$. Namely, $Q\lesssim A$.
@@
      Inclusiveness: $A$ is inclusive in $T$. Namely, $\mbox{mng}(A)\subseteq\mbox{mng}(T)$.
@@ 2.3 Fault-avoidance LLM-OM
  • $M$ is consistent if, for any user inquiries $Q$ and $Q^{\prime}$ posed at different points in time, if $\mbox{mng}(Q^{\prime})=\mbox{mng}(Q)$, then $M$ outputs an answer $A$ to $Q$ and an answer $A^{\prime}$ to $Q^{\prime}$ such that $\mbox{mng}(A)=\mbox{mng}(A^{\prime})$.
  • Suppose $M$ is augmented with input and output being $(T,Q)$ and $A$. We say that $A$ is correct for $Q$ with respect to $T$ if the following conditions hold:
@@
      Compliance: $A$ complies with the ground truth $T$ with respect to $Q$. Namely, $A\lesssim_{Q}T$.
@@
      Exclusiveness: $A$ doesn’t include unnecessary information outside the scope of $Q$. In other words, if we let $\mbox{intent}(Q)$ denote the set of intended meanings of $Q$, then $\mbox{mng}(A)\subseteq\mbox{intent}(Q)$.
@@
      Completeness: $A$ doesn’t miss any information contained in $T$ that are relevant to $Q$. In other words, $(\mbox{mng}(T)-\mbox{mng}(A))\cap\mbox{intent}(Q)=\emptyset$.
    4. 4.
      Distribution agreement: If $A$ has more than one meaning, then $\mbox{dist}(A)=\mbox{dist}(T_{A})$, where $\mbox{dist}(A)$ denotes the distribution of meanings within answer $A$ and $\mbox{dist}(T_{A})$ denotes the distribution of meanings in the corresponding text for $A$ in the augmented text $T$.
    We say that $M$ is correct with respect to $T$ if for any input $(T,Q)$, $M$ returns an answer $A$ that is correct for $Q$ with respect to $T$.
@@
      Non-exclusiveness: $A$ may contain information that extend beyond the specific information requested in the user inquiry $Q$. Namely, $\mbox{mng}(A)-\mbox{intent}(Q)\not=\emptyset$.
@@
      Incompleteness: $A$ may miss information from $T$ that pertains to $Q$. Namely, $(\mbox{mng}(T)-\mbox{mng}(A))\cap\mbox{intent}(Q)\not=\emptyset$.
@@
      Distribution discrepancy: $\mbox{dist}(A)\not=\mbox{dist}(T_{A})$.
    Essentially, adequacy acknowledges that the answer $A$ may not be perfect, but it provides valuable information to the query from the augmented text.
    We say that $M$ is adequate if for all inputs $(T,Q)$, $M$ returns an answer $A$ that is adequate for $Q$ with respect to $T$.
@@
Verifying the correctness or adequacy of the answer $A$ for a given query $Q$ with respect to $T$ calls for concrete implementations of $\mbox{toc}(X)$, $\mbox{mng}(X)$, $\mbox{intent}(Q)$, and $\mbox{dist}(X)$ using techniques in natural language processing, machine learning, deep learning, and other methods.
Consistence, however, is harder to verify. We may aim to develop a method that provides a certain guarantee that $M$ is consistent with a desired high probability.